UT 20131024 LUCI2 Daytime Commissioning

Observer: PBuschkamp, Wseifert (both remote)
Telescope Operator: NA
LBTO Support: NA

Summary

Due to urgent installation of rigid secondary at SX due to failure of AdSec
only a couple of hours for tests were available.
Walter did some basic function tests, Peter tests with the MOS
There is an issue with the Rack Cooling Control

Preparations

Scripts for CW/FW/Grating-pos/til repro are prepared and in folder daytime_com
Script for detector focus also prepared by WSeifert

Details

ROT=299deg
Basic function tests:
- FW1 & 2: ok
- Grating Sele: ok
- Grating tilt: ok
G210 -0.34/+4.35/-4.33 +4.0/-4.0/0.0 ok
G200 +1.80/+5.00/-3.78 +4.5/-3.5/0.0 ok
G150 +1.70/+3.82/-3.35 +3.5/-3.0/0.0 ok
- Image N375, SpectroSieve #0001
- Focus: neglim not working reliably here!
-> leave poslim as reference
- best focus (manually determined)
N3.75 -370 (@ T=80K) #0002 Pos 1534,1814
N1.80 -310 #0003 Pos (upper right pinhole) 1268,1397

- Peter MOS Test @210deg

- Issue with Rack Cooling Control -> Micheal, Alexander + Jenny iterating on

-- JochenHeidt - 25 Oct 2013
Topic revision: r7 - 25 Oct 2013, JochenHeidt
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback